# 22mm22mm25mm25mm

### v1

@@ 4. Simultaneous block diagonalization of Hermitian matrices
[…]
Due to their similar algebraic structures, we may apply center $Z(A_{1},A_{2}$, $\ldots,A_{m})$ in Equation (4.1) to simultaneously block diagonalize a set of Hermitian matrices via *-congruence.
@@
    There is a one-to-one correspondence between simultaneous block diagonalizations of $A_{1},A_{2},\ldots,A_{m}$ via *-congruence and complete sets of orthogonal idempotents of $Z(A_{1},A_{2},\ldots,A_{m})$.
@@
    $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalizable via *-congruence if and only if $Z(A_{1}$, $A_{2},\ldots,A_{m})$ has a complete set of $n$ orthogonal idempotents.
@@
The center of a set of Hermitian matrices in Definition 4.2 uses a conjugate transpose instead of the transpose used by the center of a set of symmetric matrices in Definition 2.1. Since both of them obey the reverse-order law, one can easily derive a proof for item (1) based on the proof in Theorems 2.7. Note that $X\in Z(A_{1},A_{2},\ldots,A_{m})$, but $iX$ may not $\in Z(A_{1},A_{2},\ldots,A_{m}),$ so $Z(A_{1},A_{2},\ldots,A_{m})$ is only a real subspace of $\mathbbm{C}^{n\times n}$. Thus, $(Z(A_{1},A_{2},\ldots,A_{m}),\odot)$ is a Jordan subalgebra of ${(\mathbbm{C}^{n\times n})}^{+}$ both viewed over the reals. Similarly, proof for items (2) to (4) can be reproduced based on the proof for Theorems 2.8. ∎
The SBDC of a set of Hermitian matrices follows the same procedure as the SBDC of a set of symmetric matrices. Thus, we adopt Algorithm 1 for the SBDC of a set of Hermitian matrices and provide two examples to demonstrate the process. Example 4.4 shows the SDC of three Hermitian matrices and Example 4.5 illustrates the SBDC of two Hermitian matrices.
@@
We then compute a pair of orthogonal idempotents based on $Z(A_{1},A_{2},A_{3})$ as
@@
We can compute their center as
$Z(A_{1},A_{2})=\left\{\begin{pmatrix}a+8b&6bi&-6bi\\ 3bi&a-b&3b\\ -2bi&2b&a\end{pmatrix}\vline\,a,b\in\mathbb{R}\right\}$
and then obtain a pair of orthogonal idempotents
@@
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}a+b-bi&-bi\\ -b&a+bi\end{pmatrix}\vline\,a,b\in\mathbb{R}\right\}.$
@@
Since such $a,b\notin\mathbb{R}$, a nontrivial idempotent does not exist. Indeed, this can also be seen by the fact that $Z(B_{1},B_{2})\cong\mathbbm{C}$ as algebras. Thus, $B_{1},B_{2}$ are not simultaneously diagonalizable via *-congruence and $A_{1},A_{2}$ are simultaneously block diagonalizable via *-congruence.

### v2

@@ 4 Simultaneous block diagonalization of Hermitian matrices
$Z(A_{1},A_{2},\dots,A_{m}):=\{X\in\mathbbm{C}^{n\times n}\mid(A_{i}X)^{*}=A_{i}X,\ 1\leq i\leq m\}.$ (4.1)
It is clear that the center of a set of Hermitian matrices is a linear space over the real field $\mathbbm{R}$ and contains all scalar matrices. Due to their similar algebraic structures, we may apply center $Z(A_{1},A_{2},\ldots,A_{m})$ in Equation (4.1) to simultaneously block diagonalize a set of Hermitian matrices via *-congruence.
@@
    There is a one-to-one correspondence between simultaneous block diagonalizations of $A_{1},A_{2},\ldots,A_{m}$ via (unitary) *-congruence and complete sets of (Hermitian) orthogonal idempotent elements of $Z(A_{1},A_{2},\ldots,A_{m})$.
@@
    $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalizable via *-congruence if and only if $Z(A_{1},A_{2},\ldots,A_{m})$ has a complete set of $n$ orthogonal idempotent elements.
  • (5)
    The finest simultaneous block diagonalization for $A_{1},A_{2},\ldots,A_{m}$ via *-congruence is unique up to equivalence.
@@
Similar to Theorem 2.4 and Corollary 2.5. ∎
The SBDC of a set of Hermitian matrices follows the same procedure as the SBDC of a set of symmetric matrices. Thus, we adopt Algorithm 1 for the SBDC of a set of Hermitian matrices and provide two examples to illustrate the method. Example 4.4 shows the SDC of three Hermitian matrices, and Example 4.5 illustrates the SBDC of two Hermitian matrices. The unitary SBDC of three Hermitian matrices is given in Example 4.6.
@@
For unitary *-congruence, we require $A_{1},A_{2},A_{3}$ to commute with each other. A direct calculation gives
$A_{1}A_{2}=\begin{pmatrix}4&9+9i\\ 8-8i&8\end{pmatrix},\quad A_{2}A_{1}=\begin{pmatrix}4&4+4i\\ 9-9i&11\end{pmatrix}$.
Since $A_{1}A_{2}\neq A_{2}A_{1}$, they do not commute and thus $A_{1},A_{2},A_{3}$ are not simultaneously diagonalizable via unitary *-congruence. We then compute a pair of orthogonal idempotents for general congruence based on $Z(A_{1},A_{2},A_{3})$ as
@@
We compute their center as
$Z(A_{1},A_{2})=\left\{\begin{pmatrix}a+8b&6bi&-6bi\\ 3bi&a-b&3b\\ -2bi&2b&a\end{pmatrix}\vline\,a,b\in\mathbbm{R}\right\}.$
By imposing the Hermitian condition on $Z(A_{1},A_{2})$, the only solution is $X=\lambda I$ for some $\lambda\in\mathbbm{C}$. Thus, $A_{1}$ and $A_{2}$ are not simultaneously block diagonalizable via unitary *-congruence. We then obtain a pair of orthogonal idempotents
@@
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}a+b-bi&-bi\\ -b&a+bi\end{pmatrix}\vline\,a,b\in\mathbbm{R}\right\}.$
@@
Since such $a,b\notin\mathbb{R}$, a nontrivial idempotent does not exist. Indeed, this can also be seen by the fact that $Z(B_{1},B_{2})\cong\mathbbm{C}$ as $\mathbbm{R}$-algebras. Thus, $B_{1},B_{2}$ are not simultaneously diagonalizable via *-congruence and $A_{1},A_{2}$ are simultaneously block diagonalizable via *-congruence.
Example 4.6.
Consider the following Hermitian matrices
$A_{1}=\begin{pmatrix}9&-6i&3i\\ 6i&9&-6i\\ -3i&6i&9\end{pmatrix},\quad A_{2}=\begin{pmatrix}-4&-2&5\\ -2&8&-2\\ 5&-2&-4\end{pmatrix},\quad A_{3}=\begin{pmatrix}15&-6&6\\ -6&6&12\\ 6&12&6\end{pmatrix}$.
First, we compute their center as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a+4b&2b&4b\\ 2b&a+b&2b\\ 4b&2b&a+4b\end{pmatrix}\vline\,a,b\in\mathbbm{R}\right\}$
and obtain a pair of Hermitian orthogonal idempotents
$\epsilon_{1}=\frac{1}{9}\begin{pmatrix}4&2&4\\ 2&1&2\\ 4&2&4\end{pmatrix}\quad\epsilon_{2}=\frac{1}{9}\begin{pmatrix}5&-2&-4\\ -2&8&-2\\ -4&-2&5\end{pmatrix}$.
This yields an invertible matrix $P=\frac{1}{3}\begin{pmatrix}2&2&i\\ 1&-2&2i\\ 2&-1&-2i\end{pmatrix}$ such that
We can now simultaneously block diagonalize $A_{1},A_{2},A_{3}$ via unitary *-congruence as
$P^{*}A_{1}P=\begin{pmatrix}9&0&0\\ 0&9&9\\ 0&9&9\end{pmatrix},\quad P^{*}A_{2}P=\begin{pmatrix}0&0&0\\ 0&0&-9i\\ 0&9i&0\end{pmatrix},\quad P^{*}A_{3}P=\begin{pmatrix}18&0&0\\ 0&18&0\\ 0&0&-9\end{pmatrix}$.
We then verify if their bottom right submatrices
$B_{1}=\begin{pmatrix}9&9\\ 9&9\end{pmatrix},\quad B_{2}=\begin{pmatrix}0&-9i\\ 9i&0\end{pmatrix},\quad B_{3}=\begin{pmatrix}18&0\\ 0&-9\end{pmatrix}$
can be simultaneously diagonalized via *-congruence. Their center is calculated as
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}a&0\\ 0&a\end{pmatrix}\vline\,a\in\mathbbm{R}\right\},$
which cannot produce nontrivial idempotent elements. Therefore, $B_{1},B_{2},B_{3}$ are not simultaneously diagonalizable via congruence, and the unitary SBDC of $A_{1},A_{2},A_{3}$ is the finest decomposition.
Use of AI tools declaration
The authors declare they have not used Artificial Intelligence (AI) tools in the creation of this article.
Acknowledgements
Conflict of interest
The authors declare no conflicts of interest.